# Superconducting Diode Effect in Weak Localization Regime

### v1

@@ Abstract
We study a dirty two-dimensional superconductor with Rashba spin-orbit coupling and in-plane Zeeman fields described by the nonlinear sigma model that includes short-range electron-electron interactions from the Coulomb and Cooper channels. The renormalized Ginzburg-Landau theory, which includes the weak localization effects at the one-loop level, is constructed by using the Keldysh functional formalism. It is shown that the tricritical point appears in the phase diagram. The superconducting diode quality factor increases divergently as the system approaches the tricritical point. Near the superconducting phase transition lines, the absolute value of the diode quality factor decreases due to the cooperation of localization and interactions. The normal conductivity of the resistive state, in which the superconducting state is destroyed by the critical current, is calculated, and localization behaviors are demonstrated.
@@
Introduction: The phenomenon in which the critical currents of superconductors become nonreciprocal, known as the superconducting diode (SD) effect [1, 2], has attracted much attention. Theoretical studies have been carried out for various types of noncentrosymmetric superconductors [3, 4, 5, 6, 7, 8, 9]. Not only clean systems but also systems with uncorrelated impurities have been studied in detail at the mean-field level [10]. If electron systems are disordered enough, they are well described by the nonlinear sigma model (NLSM) [11, 12, 13, 14], and dirty superconductors can also be described by the NLSM [15, 16, 17, 18]. Recently, an NLSM for a dirty superconductor with spin-orbit coupling (SOC) and a Zeeman field was derived [19] and used to investigate the SD effect in Josephson junctions at the mean-field level [20]. However, the Anderson localization effect, a correction to the mean-field theory in dirty systems, has not yet been highlighted in studies of the SD effect.
@@
The role of Anderson localization in superconductors is an important topic of ongoing debate. Without considering strong mesoscopic correlations and electron-electron (e-e) interactions, the properties of dirty $s$-wave superconductors at zero magnetic field remain unchanged with respect to the strength of non-magnetic disorder in the weak localization (WL) regime, known as the Anderson theorem [21]. If the Coulomb interaction is concerned, the transition temperature for the $s$-wave pairing states is suppressed in the WL regime [22, 23, 24]. The opposite effect can occur in the strong localization regime by the multifractality effect [25, 26, 27, 28]. It was also shown that, in the WL regime, the Coulomb interaction suppresses the critical current of conventional dirty superconductors at zero magnetic field and low temperatures [29, 30].
@@
In this Letter, we study the interplay of the e-e interaction and disorder on the SD effect and clarify the WL correction to conductivity when the superconducting (SC) state is broken by the critical current. We work with the NLSM for a dirty two-dimensional (2D) noncentrosymmetric superconductor, including the interactions from the Coulomb and Cooper channels in the short-range limit, and establish a renormalized Ginzburg-Landau (GL) theory at the one-loop level. While many studies [28, 31, 32, 33] have been conducted in the replica formalism, we formulate the theory with the Keldysh functional formalism [14, 15, 17, 34].
@@
Model: We consider a dirty 2D superconductor aligned in the $xy$-plane with the Rashba SOC and a Zeeman field in the $x$-direction. In the Keldysh functional formalism, the partition function is given as $Z=\int D[Q,\Delta]\exp\left({iS[Q,\Delta]}\right)$ with [19]
$S[\Delta,Q]$ $=S_{\rm HS}[\Delta]+S_{\rm N}[\Delta,Q]+S_{\rm T}[Q],$ (1)
where $S_{\rm HS}$ is responsible for the Hubbard-Stratonovich transformation associated with the SC state, $S_{\rm N}$ is the conventional non-interacting part, and $S_{\rm T}$ is the topological part. Their explicit forms are shown in Supplemental material [35]. A $\times 8$ 8 matrix field $Q$ satisfies the conditions $Q^{2}=1$ and $\mathrm{Tr}\,\{Q\}=0$. Let $\gamma_{\mu},~\tau_{\mu}$, and $\sigma_{\mu}~(\mu=0,\dots,4)$ be the Pauli matrices in the Keldysh, Nambu, and spin spaces, respectively. In particular, we write $\gamma_{\rm c}=\gamma_{0}$ and $\gamma_{\rm q}=\gamma_{1}$ and often abbreviate $\gamma_{0}$ for simplicity. The order parameter of superconductivity, that is, the SC gap, is then described by the matrix $\gamma_{\rm c}\check{\Delta}_{\rm c}+\gamma_{\rm q}\check{\Delta}_{\rm q}$, where $\check{\Delta}_{\rm c(q)}=\tau_{+}\Delta_{\rm c(q)}+\tau_{-}\Delta^{*}_{\rm c(q)}$ with $\tau_{\pm}=(\tau_{1}\pm i\tau_{2})/2$ and $\Delta_{\rm c(q)}$ is the classical (quantum) component of the SC gap. The check symbol $\check{(\cdot)}$ highlights that the objects are $\times 4$ 4 matrices. The terms in the form $\gamma_{\mu}\check{A}$ stand for the expression $\gamma_{\mu}\otimes\check{A}$, where $\otimes$ is a tensor product. We assume that the SC state is a helical state modulating in the $y$-direction and time-independent. In this case, the Cooper pair momentum corresponds to a uniform gauge field along the $y$-axis $A_{y}$, and we set $A_{x}=0$ from now on.
In addition to Eq. (1), we take into account interaction terms of the Cooper and Coulomb channels. The former can be obtained directly by integrating the spatial fluctuation of the gap function from Eq. (1) [18, 28]. The procedure results in two terms in action. The first term $\bar{S}_{\rm HS}$ has the same form as $S_{\rm HS}$ [Eq. (S.1) in Supplemental Material [35]], but the spatial modulation in the gap functions is removed. Thus, in the following, $\Delta_{\rm c(q)}$ denotes the spatially uniform gap function. The second term is
$S^{\rm(c)}_{\rm int}[Q]=\tilde{\Gamma}_{\rm c}\!\!\sum\limits_{i=1,2}\!\int_{t,\bm{r}}\!\!\!\mathrm{tr}\!\left\{{\gamma_{\rm c}\tau_{i}\delta Q_{tt}(\bm{r})}\right\}\!\mathrm{tr}\!\left\{{\gamma_{\rm q}\tau_{i}\delta Q_{tt}(\bm{r})}\right\}\!,$ (2)
@@
$S^{\rm(s)}_{\rm int}[Q]=\tilde{\Gamma}_{\rm s}\int_{t,\bm{r}}\mathrm{tr}\left\{{\gamma_{\rm c}\tau_{0}Q_{tt}(\bm{r})}\right\}\mathrm{tr}\left\{{\gamma_{\rm q}\tau_{0}Q_{tt}(\bm{r})}\right\},$ (3)
where $\tilde{\Gamma}_{\rm s}=-\pi^{2}\nu\Gamma_{\rm s}/16$ and $\Gamma_{\rm s}=1$ in the short-range limit. Thus, the total action in the partition function $Z$ is given by $S_{\rm tot}=\bar{S}_{\rm HS}+S_{\rm N}+S_{\rm T}+S^{\rm(c)}_{\rm int}+S^{\rm(s)}_{\rm int}$.
Effective action: In the ordinary mean-field theory, the system is described by saddle points of action, which are obtained by searching the matrix field $Q=\bar{Q}$ that optimizes the action under the constraints $\bar{Q}^{2}=1$ and $\mathrm{Tr}\,\{\bar{Q}\}=0$. Going beyond the mean-field approximation can be done by integrating out the fluctuations around the saddle points, resulting in the effective action of the gap function. In this work, we consider the Gaussian fluctuations around a SC saddle point $\bar{Q}_{\rm s}$. Since $\bar{Q}_{\rm s}$ must be on the same manifold as a normal saddle point $\bar{Q}_{\rm N}$, there is a matrix $U$ such that $\bar{Q}_{\rm s}=U\bar{Q}_{\rm N}U^{-1}$. The saddle point configuration may be taken as the fermionic action formulated by Keldysh path integration [14]. That is $\bar{Q}=U_{\rm K}\Lambda U_{\rm K}$, where $\Lambda=\mathrm{diag}\,\{\check{\Lambda}_{\rm R},\check{\Lambda}_{\rm A}\}$ and $\rm R(A)$ denotes the retarded (advanced) component. In the energy representation, $U_{\rm K}$ is
@@
Next, we perform the transformation $Q=U\tilde{Q}U^{-1}$ and then the transformation $\tilde{Q}=\exp\left({W/2}\right)\Lambda_{\rm N}\exp\left({-W/2}\right)$. Thus, $Q$ can be expanded to the quadratic order of $W$ as $Q=\Lambda_{\rm N}+\frac{1}{2}\,[W,\Lambda_{\rm N}]+\frac{1}{8}\,[W,[W,\Lambda_{\rm N}]]$. Expanding the total action to the quadratic order of $W$ results in $S_{\mathrm{tot}}[\Delta,Q]\rightarrow S_{\mathrm{tot}}[\Delta,\theta,W]$, where
@@
Here, $S_{0}[W]$ is the action of the free cooperon and diffuson modes, which are determined from the condition $\{W,\Lambda_{\rm N}\}=0$. The explicit expressions for each term except the last term in Eq. (5) are listed in [35]. The last term $S_{\rm II}$ contains the terms in the form $\sim\mathrm{Tr}\,\{(\dots)WW\}$ and $\sim\mathrm{Tr}\,\{W(\dots)W\}$, which do not make any contribution in the first order of $\Delta_{\rm q}$ and thus are neglected.
Now, we can find the effective action of the gap function, $S_{\rm eff}$. Regarding the terms other than $S_{0}[W]$ as perturbative terms, $S_{\rm eff}$ is given as $\exp\left({iS_{\rm eff}}\right)=\braket{\exp\left({i\left[{S_{\rm MF}+S_{\rm int}+S_{\rm I}}\right]}\right)}_{0}$, where $\braket{\cdots}_{0}$ denotes the average with respect to $S_{0}$. Performing the cumulant expansion, at the one-loop level, the effective action is obtained from the following equation:
@@
SC saddle point: The SC saddle points $\bar{Q}_{\rm s}$ are usually obtained by numerical techniques. However, since we aim to integrate out the fluctuations to obtain an effective action of the SC gap function, saddle points can be approximated so that all the fluctuations can be considered. Although taking into account all fluctuations is a difficult task, a method can give the familiar GL action for conventional dirty superconductors by calculations at the tree level [16, 17]. Here, we choose the time-independent SC saddle point $\bar{Q}_{\rm s}$ that includes the Zeeman coupling but excludes the interaction terms, the SOC, and the Cooper pair momentum. The saddle point $\bar{Q}_{\rm s}$ is obtained by solving the equation $[\check{E}_{\epsilon}+\check{\Delta}_{\rm c}+\gamma_{\rm q}\check{\Delta}_{\rm q},\bar{Q}_{\mathrm{s},\epsilon\epsilon}]=0$, where $\check{E}_{\epsilon}=\tau_{3}(\epsilon+h\sigma_{3})$ and $h$ is a Zeeman field. The equation can be easily solved if $\Delta_{\rm q}=0$. However, in the absence of the quantum part $\Delta_{\rm q}$, the solution cannot include any effect of the terms other than the term associated with $\check{E}$ on the effective action. We overcome this problem by instead using the solution of the following equation:
$[\check{E}_{\epsilon}+\check{\Delta}_{\mathrm{R(A)},\epsilon}$ $,\check{\Lambda}_{\mathrm{sR(A)},\epsilon\epsilon}]=0,$ (7)
$\check{\Delta}_{\mathrm{R(A)},\epsilon}=$ $\check{\Delta}_{\rm c}\pm F_{\epsilon}\check{\Delta}_{\rm q}.$ (8)
@@
and $\epsilon_{\pm}=\epsilon\pm h$. $\hat{X}_{\rm A}$ has the same form as $\hat{X}_{\rm R}$ but $\theta_{\mathrm{R},\epsilon\pm}$ is replaced by $\theta_{\mathrm{A},\epsilon\pm}$, which satisfies $\tanh\theta_{\mathrm{A},\epsilon\pm}=\Delta_{\mathrm{A},\epsilon}/|\epsilon_{\pm}\!-\!i0|$. This is indeed a possible form of the spectral angle mentioned below Eq. (4). Defining $U\equiv U_{\rm K}U_{\rm s}$ and $U_{\rm s}\equiv\mathrm{diag}\left\{{\check{U}_{\rm sR},\check{U}_{\rm sA}}\right\}$, we can rewritte $\bar{Q}_{\rm s}$ as $\bar{Q}_{\rm s}=U\Lambda_{\rm N}U^{-1}$.
The solution of Eq. (7) can be used to obtain the effective action including the contributions of the SOC, Cooper-pair momentum, and interactions. However, we will go to a more accurate approximation by allowing the spectral angles $\theta_{\mathrm{R(A)},\epsilon\pm}$ to be determined by optimizing the total effective action with respect to them. The outcome is the so-called modified Usadel equation, first proposed in Ref. [28] by the replica formalism. To derive a similar equation in the Keldysh functional formalism, we have to take into account the quantum part of the spectral angle. Otherwise, for the same reason as described in the previous paragraph, the effective action becomes trivial. The problem is how the quantum part of the spectral angles should enter $\Lambda_{\rm s}$. By observing Eqs. (8) and (10), we arrive at the ansatz that the generalized spectral angles should have the form like Eq. (8) for the gap function, and the spectral angles in Eq. (9) should be replaced as follows:
$\theta_{\mathrm{R},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{R},\epsilon\pm}=\theta^{\rm c}_{\mathrm{R},\epsilon\pm}\!\!+F_{\epsilon}\theta^{\rm q}_{\mathrm{R},\epsilon\pm},$ (11)
$\theta_{\mathrm{A},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{A},\epsilon\pm}=\theta^{\rm c}_{\mathrm{A},\epsilon\pm}\!\!-F_{\epsilon}\theta^{\rm q}_{\mathrm{A},\epsilon\pm},$ (12)
@@
where $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$ represent the classical and quantum parts of the generalized spectral angles, respectively. In the above equation, $F_{\epsilon}$ plays an essential role. The effective action would become trivial if there were no term that couples $F_{\epsilon}$ to $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$. Hereafter, we adopt the SC saddle point $\bar{Q}_{\rm s}$ that is determined by Eq. (9) with Eqs. (11) and (12). The quantum parts will be set to zero after performing the first derivatives with respect to them.
@@
Modified Usadel equation: Following the plan described in the previous paragraph, we determine the form of $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}$ by optimizing the effective action $S_{\rm eff}$ with respect to each spectral angle. We work within GL theory by expanding $\theta_{\mathrm{R(A)},\epsilon\pm}$ as $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}=\Delta_{\rm c}f^{\rm R(A)}_{1\epsilon\pm}+\Delta^{3}_{\rm c}f^{\rm R(A)}_{3\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}=\theta^{\rm R(A)q}_{1\epsilon\pm}+\theta^{\rm R(A)q}_{3\epsilon\pm}$. In this way, $f^{\rm R(A)}_{j\epsilon\pm}$ is determined by the following equation:
$[\delta S_{\rm{eff}}/\delta\theta^{\mu\mathrm{q}}_{j\epsilon\pm}]_{\theta^{\mathrm{q}}\rightarrow 0}=0,$ (13)
@@
where $\mu=\rm R,A$ and $j=1$ or 333. In the following, $f^{\rm R(A)}_{j\epsilon\pm}$ is referred to as spectral function. In the calculation of Eq. (13), the retarded and advanced parts are not independent. Otherwise, irregular divergences appear in the left-hand side of Eq. (13). This argument is explained in detail in Supplemental Material [35]. We find that all irregular terms vanish if the following condition holds:
@@
Using Eqs. (11)--(14) and performing analytic continuation, we obtain the modified Usadel equations at the first and third orders of $\Delta_{\rm c}$ as follows:
@@
where $f_{jn\pm}=if^{\rm R}_{ji\omega_{n}\pm}$, $\omega_{n\pm}=\omega_{n}\pm ih$, $\omega_{n}=(2n+1)\pi T~(n\in\mathbb{Z})$, and $\mathcal{S}_{\mathrm{source},1}=1$. The others are shown in Supplemental Material [35] to avoid describing the lengthy equations, and we find that the Cooper interaction terms are not involved in the equation determining $f_{1n\pm}$. The equations are solved numerically by the Padé decomposition method [36], which can achieve fast convergence when calculating the Matsubara summation.
Equation of state: The gap equation is obtained from the stationary condition $[\partial S/\partial\Delta_{\rm q}]_{\Delta_{\rm q}\rightarrow 0}=0$. In GL theory, the equation has the form $\alpha_{2}\Delta_{\rm c}+\alpha_{4}\Delta_{\rm c}^{3}=0$, where
$\alpha_{2}$ $=4\,\Gamma^{-1}_{\rm c}\!-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big{[}f_{1n\lambda}\!+\!\frac{1}{\omega_{n}\!+\!i\lambda h}\Big{]}$

### v2

@@ Abstract
We study a dirty two-dimensional superconductor with Rashba spin-orbit coupling and in-plane Zeeman fields described by the nonlinear sigma model that includes the Cooper and long-range Coulomb interactions. The renormalized Ginzburg-Landau theory, which includes the weak localization effects at the one-loop level, is constructed using the Keldysh functional formalism. It is shown that the transition temperature and magnetic field, as well as the tricritical point appearing in the phase diagram, are suppressed by the interactions. Nevertheless, we have found a universal behavior in the high-transition-temperature regime that demonstrates the robustness of the superconducting diode effect against the interactions. The conductivity of the resistive states emerging after the superconducting states are destroyed by the critical current is also calculated, and localization behaviors are demonstrated.
@@
Introduction: The phenomenon in which the critical currents of superconductors become nonreciprocal, known as the superconducting diode (SD) effect [1, 2], has attracted much attention. Theoretical studies have been carried out for various types of noncentrosymmetric superconductors [3, 4, 5, 6, 7, 8, 9]. Not only clean systems but also disordered systems have been studied at the mean-field (MF) level [10, 11, 12]. If electron systems are disordered enough, they are well described by the nonlinear sigma model (NLSM) [13, 14, 15, 16], and dirty superconductors can also be described by the NLSM [17, 18, 19, 20]. Recently, an NLSM for a dirty superconductor with spin-orbit coupling (SOC) and Zeeman field was derived [21] and used to investigate the SD effect in Josephson junctions at the MF level [22]. However, the Anderson localization effect, a correction to the MF theory in dirty systems, has not yet been highlighted in studies of the SD effect.
@@
The role of Anderson localization in superconductors is an important topic of ongoing debate. Without electron-electron (e-e) interactions, the properties of dirty $s$-wave superconductors at zero magnetic field remain unchanged with respect to the strength of non-magnetic disorder in the weak localization (WL) regime, known as the Anderson theorem [23]. However, if the long-range Coulomb interaction exists, the transition temperature is largely suppressed even in the WL regime [24, 25, 26]. Moreover, the transition temperature can also be enhanced by the multifractality effect [27, 28, 29, 30], where the short-range nature of the e-e interactions is important. It was also shown that, in the WL regime, the Coulomb interaction lowers the critical current of conventional dirty superconductors at zero magnetic field and low temperatures [31, 32].
@@
In this Letter, we study the role of WL correction from the Cooper and long-range Coulomb interactions, which are ubiquitous in the solid state systems [33, 34, 35, 36, 37], on the SD effect. In normal metals, the localization behavior can be changed by the SOC. It is interesting to consider whether the WL behavior of the resistive states that emerge after the superconducting (SC) states are destroyed by the critical currents can be enhanced while maintaining the high efficiency of the SD effect. We thus also investigate the relationship between the localization behavior in such resistive states and the SD effect. We work with the NLSM for a dirty two-dimensional (2D) noncentrosymmetric superconductor, and establish a renormalized Ginzburg-Landau (GL) theory at the one-loop level. While many studies [30, 38, 39, 40] have been conducted in the replica formalism, we formulate the theory with the Keldysh functional formalism [16, 17, 19, 41].
@@
Model: We consider a dirty 2D superconductor aligned in the $xy$-plane with the Rashba SOC and a Zeeman field $h$ in the $x$-direction. It was shown in [21] that in the Keldysh functional formalism, the partition function of the system without e-e interactions is given as $Z=\int D[Q,\Delta]\exp\left({iS[Q,\Delta]}\right)$, where
$S[\Delta,Q]$ $=S_{\rm HS}[\Delta]+S_{\rm N}[\Delta,Q]+S_{\rm T}[Q].$ (1)
@@
In Eq. (1), the fluctuation of the gap matrix $\Delta$ from the static helical SC state with a fixed Cooper pair momentum $\bm{P}$ is already integrated out. Here, the static helical SC state is characterized by the gap function $\Delta_{\rm cl(q)}\exp\,(i\bm{P}\cdot\bm{r})$, where $\Delta_{\rm cl(q)}\in\mathbb{R}$ is the classical (quantum) component of the SC gap, and it does not depend on space and time. See the procedure in [42]. Thus, only the field $Q$ and the static uniform SC gaps remain in the integral measure of the partition function $Z$. Moreover, since we consider the Zeeman field $h$ pointing in the $x$-direction, the SD effect occurs when a supercurrent flows in the $y$-direction. We thus set $P_{x}=0$. The factor $\exp\,(iP_{y}y)$ can also be cancelled out by a U(1) transformation. As a result, in Eq. (1), a uniform gauge field, $A_{y}=P_{y}/2$, appears in the covariant derivative, and $\Delta$ is expressed as $\Delta=\gamma_{\rm cl}\check{\Delta}_{\rm cl}+\gamma_{\rm q}\check{\Delta}_{\rm q}$, where $\check{\Delta}_{\rm cl(q)}=i\tau_{2}\Delta_{\rm cl(q)}$.
In addition to Eq. (1), we take into account interaction terms of the Cooper and Coulomb channels. Let $S^{\rm(c)}_{\rm int}$ and $S^{\rm(s)}_{\rm int}$ denote the Cooper and Coulomb interaction actions, respectively. Their explicit forms are shown in [42]. Note that we consider the Coulomb interaction in the long-range limit. The total partition function $Z_{\rm tot}$ is given as $Z_{\rm tot}=\int D[Q,\Delta]\exp\left({iS_{\rm tot}[Q,\Delta]}\right)$, where
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}S_{\rm tot}[\Delta,Q]=S[\Delta,Q]+S^{\rm(c)}_{\rm int}[Q]+S^{\rm(s)}_{\rm int}[Q].$} (2)
Despite being beyond our scope, it is instructive to note that the triplet channel of e-e interactions can be neglected in any case if the triplet diffusive modes acquire a large mass. The diffusive modes, which are cooperon and diffuson modes, will be described later. In our case, the mass of the triplet diffusive modes is the DP relaxation rate $\Gamma_{\rm r}$. Further details are discussed in [42].
Effective action: In the NLSM, the system is described at the MF level by searching the saddle matrix field $Q=\bar{Q}$ that optimizes the action under the constraints $\bar{Q}^{2}=1$ and $\mathrm{Tr}\,\{\bar{Q}\}=0$. It is possible to optimize the total action $S_{\rm tot}$ to obtain an MF theory that includes the contribution of e-e interactions [30]. However, the corresponding Usadel equation has a nonlinear form that is difficult to solve. We thus take an alternative track [30]: Setting a saddle point and integrating out fluctuations around it. This ultimately leads to an effective action of the SC gap. The procedure to achieve such effective action is outlined below.
We consider the Gaussian fluctuations around an SC saddle point $\bar{Q}_{\rm s}$ whose form will be specified later. Although it is possible to adopt the normal saddle point $\bar{Q}_{\rm N}$ as the origin of expansion, the high-order calculation of perturbation theory is needed [19]. Time and effort can be saved by considering the SC saddle point [18]. Since $\bar{Q}_{\rm N}$ and $\bar{Q}_{\rm s}$ satisfy $\bar{Q}^{2}_{\rm N}=\bar{Q}^{2}_{\rm s}=1$, there is a matrix $U$ such that $\bar{Q}_{\rm s}=U\bar{Q}_{\rm N}U^{-1}$. The saddle point configuration is the same as the Keldysh fermionic action [16]. That is $\bar{Q}=U_{\rm K}\Lambda U_{\rm K}$, where $\Lambda=\mathrm{diag}\,\{\check{\Lambda}_{\rm R},\check{\Lambda}_{\rm A}\}$ and $\rm R(A)$ denotes the retarded (advanced) component. In the energy representation, $U_{\rm K}$ is
@@
Next, we perform the transformation $Q=U\tilde{Q}U^{-1}$ and then the transformation $\tilde{Q}=\exp\left({W/2}\right)\Lambda_{\rm N}\exp\left({-W/2}\right)$. Thus, $Q$ can be expanded to the quadratic order of $W$ as $\tilde{Q}=\Lambda_{\rm N}+\frac{1}{2}\,[W,\Lambda_{\rm N}]+\frac{1}{8}\,[W,[W,\Lambda_{\rm N}]]$. Expanding the total action to the quadratic order of $W$ results in $S_{\mathrm{tot}}[\Delta,Q]\rightarrow S_{\mathrm{tot}}[\Delta,\theta,W]$, where
@@
Here, $S_{0}[W]$ is the action of the free cooperon and diffuson modes, which are determined from the condition $\{W,\Lambda_{\rm N}\}=0$. The explicit expressions for each term except the last term in Eq. (4) are listed in [42]. The last term $S_{\rm II}$ contains terms in the form $\sim\mathrm{Tr}\,\{(\dots)WW\}$ and $\sim\mathrm{Tr}\,\{W(\dots)W\}$, which do not make any contribution in the first order of $\Delta_{\rm q}$ and thus are neglected.
Now, we can find the effective action of the gap function, $S_{\rm eff}$. Regarding the terms other than $S_{0}[W]$ as perturbative terms, $S_{\rm eff}$ is given as $\exp\left({iS_{\rm eff}}\right)=\braket{\exp\left({i\left[{S_{\rm MF}+S_{\rm int}+S_{\rm I}}\right]}\right)}_{0}$, where $\braket{\cdots}_{0}$ denotes the average with respect to $S_{0}$. Performing the cumulant expansion at the one-loop level, the effective action is obtained from the following equation:
@@
SC saddle point: Choosing the appropriate saddle point is a key to a good approximation of perturbation theory. It is naively expected that the best SC saddle point $\bar{Q}_{\rm s}$ should optimize the effective action Eq. (5). However, an estimation of $\bar{Q}_{\rm s}$ is necessary to make the analysis possible. This is achieved by setting a saddle point and then incorporating contributions from terms unrelated to the saddle point. We choose a saddle point that includes the Zeeman coupling. In the ordinary MF theory, this is obtained by solving equation $\check{E}_{\epsilon}+\check{\Delta}_{\rm cl}+\gamma_{\rm q}\check{\Delta}_{\rm q},\bar{Q}_{\mathrm{s},\epsilon\epsilon}]=0$ with $\Delta_{\rm q}=0$, where $\check{E}_{\epsilon}=\tau_{3}(\epsilon+h\sigma_{3})$. However, the solution does not include any effect of the terms unrelated to the saddle point because the quantum part necessary for deriving the gap equation is absent. To overcome this, we demand the retarded and advanced components of the equation to be zero, i.e.,
$\quad\quad~[\check{E}_{\epsilon}+\check{\Delta}_{\mathrm{R(A)},\epsilon},\check{\Lambda}_{\mathrm{sR(A)},\epsilon\epsilon}]=0,$ (6)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\check{\Delta}_{\mathrm{R(A)},\epsilon}=i\tau_{2}\Delta_{\mathrm{R(A)},\epsilon}\equiv i\tau_{2}\,(\Delta_{\rm cl}\pm F_{\epsilon}\Delta_{\rm q}).}$ (7)
@@
and $\epsilon_{\pm}=\epsilon\pm h$. $\hat{X}_{\rm A}$ has the same form as $\hat{X}_{\rm R}$ but $\theta_{\mathrm{R},\epsilon\pm}$ is replaced by $\theta_{\mathrm{A},\epsilon\pm}$, which satisfies $\tanh\theta_{\mathrm{A},\epsilon\pm}=\Delta_{\mathrm{A},\epsilon}/|\epsilon_{\pm}\!-\!i0|$. This is indeed a possible form of the spectral angle mentioned below Eq. (3). Defining $U\equiv U_{\rm K}U_{\rm s}$ and $U_{\rm s}\equiv\mathrm{diag}\left\{{\check{U}_{\rm sR},\check{U}_{\rm sA}}\right\}$, we can rewrite $\bar{Q}_{\rm s}$ as $\bar{Q}_{\rm s}=U\Lambda_{\rm N}U^{-1}$.
The solution presented above can be used to obtain the effective action incorporating the contributions of the SOC, Cooper-pair momentum, and interactions. This is allowed by the presence of $F_{\epsilon}$ in Eq. (7). Without this term, the effective action would become trivial through the energy integration. However, since we search the saddle point that optimizes Eq. (5), some terms in Eq. (8) need to be specified by the saddle point condition for Eq. (5). Similarly to Ref. [30], we leave the spectral angle as an unknown quality. By paying attention to the role of $F_{\epsilon}$ described above, we arrive at the ansatz that the spectral angles in Eq. (8) should be replaced by the generalized spectral angles shown below:
$\theta_{\mathrm{R},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{R},\epsilon\pm}=\theta^{\rm cl}_{\mathrm{R},\epsilon\pm}\!\!+F_{\epsilon}\theta^{\rm q}_{\mathrm{R},\epsilon\pm},$ (10)
$\theta_{\mathrm{A},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{A},\epsilon\pm}=\theta^{\rm cl}_{\mathrm{A},\epsilon\pm}\!\!-F_{\epsilon}\theta^{\rm q}_{\mathrm{A},\epsilon\pm},$ (11)
@@
where $\theta^{\rm cl}_{\mathrm{R(A)},\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$ represent the classical and quantum parts of the generalized spectral angles, respectively. Hereafter, we adopt the SC saddle point $\bar{Q}_{\rm s}$ that is determined by Eq. (8) with Eqs. (10) and (11). The quantum parts will be set to zero after performing the first derivatives with respect to them.
@@
Modified Usadel equation: Following the plan described in the previous paragraph, we determine the form of $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}$ by optimizing the effective action $S_{\rm eff}$ with respect to each spectral angle. We work within the GL theory by expanding $\theta_{\mathrm{R(A)},\epsilon\pm}$ as $\theta^{\rm cl}_{\mathrm{R(A)},\epsilon\pm}=\Delta_{\rm cl}f^{\rm R(A)}_{1\epsilon\pm}+\Delta^{3}_{\rm cl}f^{\rm R(A)}_{3\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}=\theta^{\rm R(A)q}_{1\epsilon\pm}+\theta^{\rm R(A)q}_{3\epsilon\pm}$. In this way, $f^{\rm R(A)}_{j\epsilon\pm}$ is determined by the following equation:
$\delta S_{\rm{eff}}/\delta\theta^{\mu\mathrm{q}}_{j\epsilon\pm}]_{\theta^{\mathrm{q}}\rightarrow 0}=0,$ (12)
@@
where $\mu=\rm R,A$ and $j=1$ or $3$. In the following, $f^{\rm R(A)}_{j\epsilon\pm}$ is referred to as spectral function. Note that this spectral function $f^{\rm R(A)}_{j\epsilon\pm}$ is different from the Fermi distribution function. The retarded and advanced spectral functions are not actually independent. Otherwise, irregular divergences appear in the left-hand side of Eq. (12). This argument is explained in detail in Supplemental Material [42]. We find that all irregular terms vanish if the following condition holds:
@@
Using Eqs. (10)-(13) and performing analytic continuation, we obtain the modified Usadel equation [30] in the first and third orders of $\Delta_{\rm cl}$ as follows:
@@
where $f_{jn\pm}=if^{\rm R}_{ji\omega_{n}\pm}$, $\omega_{n\pm}=\omega_{n}\pm ih$, $\omega_{n}=(2n+1)\pi T~(n\in\mathbb{Z})$, and $\mathcal{S}_{\mathrm{source},1}=1$. The others are shown in Supplemental Material [42] to avoid describing lengthy equations. We find that the correction from the long-range Coulomb interaction is much greater than that from the Cooper interaction. The equations are solved numerically by the Padé decomposition method [43], which can achieve fast convergence when calculating the Matsubara summation.
Equation of state: The gap equation is obtained from the stationary condition $\partial S/\partial\Delta_{\rm q}]_{\Delta_{\rm q}\rightarrow 0}=0$. At the GL level, the equation is in the form $\alpha_{2}\Delta_{\rm cl}+\alpha_{4}\Delta_{\rm cl}^{3}=0$, which corresponds to minimizing the GL free energy $f_{\mathrm{GL}}=\mathcal{C}\,(\frac{1}{2}\alpha_{2}\Delta^{2}_{\rm cl}+\frac{1}{4}\alpha_{4}\Delta^{4}_{\rm cl})$, where $\mathcal{C}$ is a numerical constant. The coefficients $\alpha_{2}$ and $\alpha_{4}$ are given as follows:
$\alpha_{2}$ $=4\,\Gamma^{-1}_{\rm c}\!-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big[f_{1n\lambda}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\frac{1}{\omega_{n}\!+\!i\lambda h}\Big]$